OpenStudy (anonymous):

2x-3y=4 x=-4y-9 solve using substitution

OpenStudy (anonymous):

step 1: isolate one variable in one equation

OpenStudy (anonymous):

2(-4y-9)-3y=4 -11y=22 y=-2

OpenStudy (anonymous):

give me second

OpenStudy (anonymous):

y=2 and x=-1

OpenStudy (anonymous):

@amy51275 are u sure its not -2 instead of 2? just making sure...

OpenStudy (anonymous):

-2 sorry typo

OpenStudy (anonymous):

okay just checking..thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

I need help one more time can u help @amy51275 please

OpenStudy (anonymous):

sure i can try my best to

OpenStudy (anonymous):

okay thanks... 6y=-2x+24 5x-2y=9 solve using substitution

OpenStudy (anonymous):

(3,3)